Hi there,
I am confronted with the following question:
"A job takes 4 hrs for two people working together. If one person works alone he can do the job in 6 hrs. How long will it take the other person working alone to complete the job?"
I don't really know how to approach this question. It seems so simple yet I can't seem to figure out why the answer is 12 hrs. Please help me out.
Thanks in advance,
Mike

Answer by Earlsdon(6294)

This kind of problem can be solved by asking "how much of the job is done in 1 hour?"
Let's call the two people Sam and Eric, or S and E for short.
So, S and E together can do the whole job in 4 hours. This means that they can do 1/4 of the job in 1 hour.
Now, Sam can do the whole job in 6 hours. This means that Sam can do 1/6 of the job in 1 hour.
To find find out how much of the job Eric can do in 1 hour, simply subtract what Sam can in 1 hour from what Sam & Eric can do together in 1 hour.
(S + E) - S = E
1/4 - 1/6 = 3/12 - 2/12 = 1/12
So, Eric can do 1/12 of the job in 1 hour, which means it will take him 12 hours to do the whole job.